A random sample of 16 pharmacy customers showed the waiting times below (in minutes).

Good afternoon,
Can you please help me with this problem? I don't know what I'm doing wrong, but all answers are wrong:(
Thank you
21 22 22 17 21 17 23 20
20 24 9 22 16 21 22 21


Find a 90 percent confidence interval for μ, assuming that the sample is from a normal population.
(Round your standard deviation answer to 4 decimal places and t-value to 3 decimal places. Round your answers to 3 decimal places.)

Answer by Theo(13342)

try these numbers.
see if they work.
let me know how you did.
if ok, i'll send you method.

sample mean = 19.875
sample size = 16
degrees of freedom = 15
sample standard deviation = 3.6492
critical t-score with 15 degrees of freedom at 90% two tail confidence interval = plus or minus 1.753.
t-score formula is t = (x - m) / s
x is the raw score
m is the mean
s is the standard error of the sample.
standard error of the sample = 3.6492 / sqrt(16) = .9123
on the low side, -1.753 = (x - 19.875) / /.9123
solve for x to get:
x = 18.276
on the high side, 1.753 = (x - 19.875) / .9123
solve for x to get:
x = 21.474
i used an online calculator and got:
18.27569 and 21.47431.
me and the calculator agree.
see what you get.